OpenStudy (anonymous):

9=(sqrt of 3) ^ (4x+6)... Can someone please explain their method/strategy as well? I don't understand.

OpenStudy (dumbcow):

\[(\sqrt{3})^{4x+6} = 9\] you can take log of both sides \[\log \sqrt{3})^{4x+6} = \log 9\] \[(4x+6)(\log \sqrt{3}) = \log 9\] \[4x + 6 = \frac{\log 9}{\log \sqrt{3}} = 4\]

OpenStudy (anonymous):

that is one way. without the log you can say \[\sqrt{3}^2=3\] so \[\sqrt{3}^4=9\] and therefore \[4x+6=9\] etc

OpenStudy (anonymous):

Thanks!! Both of you, yeah I needed to do it without the log.

OpenStudy (dumbcow):

oh ok, yeah mine is the formal way which will always work cause what if it was sqrt(2) then you have to use log

OpenStudy (anonymous):

Yeah, I am probably going to learn log next.. Hopefully I'll understand that.. Thanks though

OpenStudy (dumbcow):

no problem

OpenStudy (anonymous):

One more question...

OpenStudy (anonymous):

?

OpenStudy (anonymous):

If I were to explain the steps of how I did this, I'd say I created equal bases and the solved for the variable and got x=3/4

OpenStudy (dumbcow):

yes, by getting equal bases you can set the exponents equal to each other and then solve for the variable. i get x = -1/2 for this problem

OpenStudy (anonymous):

Wait how did you get -1/2

OpenStudy (dumbcow):

4x+6 = 4 4x = -2 x = -2/4 = -1/2

OpenStudy (dumbcow):

satellite had a typo at the end since 9 can be written as sqrt(3)^4 the equal base is sqrt(3) exponents are (4x+6) on left and 4 on right 4x+6 = 4